OpenStudy (anonymous):

y varies inversely with x and y = 5 when x = 2. Which is the inverse variation equation for the relationship? A.y=2x B.y=1/2x C.y=10/x D.y=x+2

OpenStudy (anonymous):

Last time I annoy you I promise! @ranga

OpenStudy (ranga):

"y varies inversely with x" means: y = k * 1/x when x =2, y = 5. Put the numbers above and solve for k.

OpenStudy (anonymous):

5=k*0.2?

OpenStudy (ranga):

x is 2 not 0.2

OpenStudy (anonymous):

oh sorry. whoops

OpenStudy (anonymous):

5=k*2?

OpenStudy (anonymous):

in other words y=2x? which would b a?

OpenStudy (ranga):

No. This is inverse variation. So y = k * 1 / x

OpenStudy (anonymous):

Then it might be B.?

OpenStudy (ranga):

Put x = 2, y = 5 in y = k * 1 / x 5 = k * 1 / 2 = k/2 multiply both sides by 2 2*5 = k k = 10 So y = 10 / x

OpenStudy (anonymous):

oh ok I get it now thanks again your a lifesaver!!!! :))))

OpenStudy (ranga):

you are welcome.